### v7

Abstract
The efficient segmentation of text information from the background in degraded color document images is an important challenge in the preservation of ancient manuscripts. The imperfect preservation of ancient manuscripts has led to various types of degradation over time, such as staining, yellowing, and ink seepage, badly affecting document image binarization results. This work proposes a three-stage method to generate binarization image results for degraded colour document images using generative adversarial networks (GANs). Stage-1 involves applying discrete wavelet transform and retaining the low-low subband images for document image enhancement. In Stage-2, the original input image is split into red, green, and blue (RGB) three single-channel images and one grayscale image, and each image is trained with independent GANs to extract color foreground information. In Stage-3, the output images of Stage-2 and the resized input images are used to train independent GANs to generate document binarization results, enabling the combination of global and local features. The experimental results show that the Avg-Score of the proposed method is 77.64, 77.95, 79.05, 76.38, 75.34, and 77.00 on the (H)-DIBCO 2011, 2013, 2014, 2016, 2017, and 2018 datasets, which achieves the state-of-the-art level. The implementation code for this work is available at https://github.com/abcpp12383/ThreeStageBinarization.
keywords:
[…]
1 Introduction
Historical analyses of human societies primarily rely on reconstructing and analyzing preserved ancient documents. Parchment manuscripts and paintings have predominantly served as the means of preserving crucial ancient documents throughout the evolution of human society. Analyzing preserved ancient documents presents challenges due to potential forms of interference and degradation, including paper yellowing, text fading, page contamination and moisture damage.
Document image enhancement is one of the most important tasks in the reconstruction and analysis of ancient documents. It involves document image binarization to extract essential text information from the page. Elevating the image quality of deteriorated documents necessitates addressing various challenges, including text degradation and bleed-through issues, such as text degradation and bleed-through [1, 2, 3, 4]. However, traditional binarization methods cannot extract text information from the background of degraded documents effectively, often resulting in incomplete removal of shadows and noises, or loss of text information [5, 6, 7].
With the development of deep learning, neural networks have been applied to image binarization. Calvo-Zaragoza and Gallego [8] introduced a selectional auto-encoder approach, parsing the text through the trained network model, and performing binarization using threshold. Fully convolutional network (FCN) [9] enables the generated images to maintain the same size as the input images, and the task [10] of generating document binarization using FCN is also included in the semantic segmentation task. However, most image binarization methods primarily process grayscale documents, targeting contaminated black and white scanned ancient documents as their default.
Considering that the images of some ancient documents are in colour, we propose a novel three-stage method for document image enhancement and colour degradation document binarization. The proposed method consists of three stages. In Stage-1, we split the original input image into red, green, and blue (RGB) three single-channel images and one grayscale image, and apply discrete wavelet transform (DWT) to retain the low-low (LL) subband images and normalize them to achieve image enhancement. In Stage-2, we train four independent adversarial networks with the four single-channel images to eliminate background information and extract color foreground details from the local image patches. During the processing of Stage-3, the multiscale adversarial networks produce $224\times 224$ image patch (local) prediction outputs and $512\times 512$ resized image (global) prediction outputs for the document images, which are subsequently fused to obtain the final output.
To evaluate the performance of the proposed method for degraded document image binarization, we compare it with other methods using DIBCO datasets of 2011, 2013, 2014, 2016, 2017, and 2018 [11, 12, 13, 14, 15, 16].
The main contributions of this paper are as follows:
  • Introducing a novel three-stage generative adversarial networks method for enhancing and binarizing the degraded color documents.
  • The incorporation of DWT prior to the training of generative adversarial networks makes the image to be input closer to the Ground Truth image.
  • The proposed method achieves the state-of-the-art performance on multiple benchmark datasets.
The rest of this paper is organized as follows: Section 2 describes the deep learning method for document image binarization. Section 3 introduces a novel three-stage generative adversarial networks method. Section 4 presents the experiments performed on this work, and compares the performance of our method and other methods on several benchmark datasets. Finally, Section 5 discusses the conclusion and the future work of this paper.
2 Related Work
In previous researches focusing on document image binarization, the field is broadly divided into two categories: traditional image processing methods and deep learning-based semantic segmentation methods. However, traditional image processing binarization methods are always based on the pixel-level local threshold values, such as global binarization [6] and local adaptive binarization [5, 7]. These methods are not satisfactory when dealing with documents affected by excessive interference or degradation [17, 18].
With the advancement of deep learning, the convolutional neural network (CNN) framework [19, 20] has become the predominant backbone of computer vision tasks. Within computer vision, semantic segmentation is one of the most important research topics, encompassing document image binarization. For instance, Tensmeyer et al. [10] employed pseudo-F-measure (p-FM) and F-measure (FM) losses during the training of FCN to apply the model to the document image binarization task. Ronneberger et al. [21] extended FCN, presenting the U-Net architecture, which captures contextual and spatial information through its distinctive U-shaped design. Inspired by this, Peng et al. [22] introduced an encoder-decoder architecture for document image binarization, where the decoder transforms the low-resolution representation to the original dimensions, resulting in the binarized image. Vo et al. [23] first applied hierarchical deep supervised network (DSN) to document image binarization and achieved the state-of-the-art (SOTA) performance on several benchmark datasets. He et al. [24] proposed an iterative deep learning framework applied to document image binarization and the trained model achieved excellent performance.
In recent years, generative adversarial networks (GANs) [25] have achieved exciting results in the generation of real-world images. Unlike GANs which generate output images from random noise, Isola et al. [26] proposed Pix2Pix GAN based on conditional GAN (cGAN) [27] for image-to-image translation. To ensure that the generated images are more realistic, Pix2Pix GAN employs generators for loss function optimization. Within the network architecture, the generator evaluates the minimized generated image against the target output image, quantifying the $L1$ loss between them. Meanwhile, the discriminator assesses the credibility of the generated image, relying on the input image and the reference target image. This process is depicted by the subsequent equation:
$\begin{split}\emph{arg}\min_{G}\max_{D}V(G,D)=\mathop{\mathbb{E}_{x}}[\log(1-D(x,G(x)))]\\ +\mathop{\mathbb{E}_{x,y}}[\log D(x,y)]+\lambda\mathop{\mathbb{E}_{x,y}}[||y-G(x)||_{1}]\end{split}$ (1)
where the input data distribution $\mathop{\mathbb{P}_{x}}$ samples the input image $x$; the ground truth image of the input image is $y$, and the hyperparameter to increase the model regularization effect is $\lambda$. Generator $G$ generates the input image as the generated image $G(x)$. From Eq. (1), the loss function between the generated image and the ground truth image is $L1$ instead of $L2$, which is beneficial to reduce blurring and ambiguity in the generation process.
Bhunia et al. [28] performed document image enhancement on the DIBCO dataset and used cGAN [27] for image binarization. Inspired by Pix2Pix GAN [26], Zhao et al. [29] proposed a cascade generator structure for document image binarization based on Pix2Pix GAN to solve the problem of multi-scale information combination. De et al. [30] designed two discriminators to combine local and global information. Suh et al. [31] set up a two-stage generative adversarial network for image binarization using Patch GAN [26]. These methods keep breaking the SOTA performance on DIBCO datasets.
Figure 1: The architecture of Stage-1 of the proposed method. The original input image is divided into patches of $224\times 224$ size, and split into grayscale images and RGB three single-channel images. The single-channel images are all in grayscale. However, to visually represent the different channels of the image separately, the single-channel images in this diagram are replaced with images in three different RGB colors.
Figure 2: The architecture of the Haar wavelet transform used by the proposed method, which takes the red single-channel image patch as an example. The processes of the transform of the green and blue single-channel image patches are the same.
3 Proposed Method
3.1 Overall Network Architecture
Considering the complexity of color document image degradation and the advantages [31] of processing different color channels separately, our proposed method divides the overall network architecture into three stages, which focus on enhancing image quality of degraded color documents and extracting text information from the images. In Stage-1, we use DWT and normalization to preprocess the patches of the input image. These preprocessed image patches are employed to train as the input of the generator within the adversarial network framework. Stage-2 focuses on the elimination of background color noise and the extraction of color foreground information, while Stage-3 involves the fusion of local and global predictions to produce the final output. The detailed introduction of each stage can be found in Sections 3.2, 3.3, and 3.4, respectively.
3.2 Stage-1
This work employs four different generators to extract foreground color information from degraded document images while eliminating the background color. To preprocess the input of the different generators, we first split the original color image (24-bit) into RGB three single-channel images (8-bit) and one grayscale image (8-bit).
The ground truth (GT) image provided by the dataset is a single-channel binary image (8-bit), where pixel points have values of 0 for characters and 1 for the background. Images with varying channel numbers cannot be directly put into the discriminator for processing. As shown in Figure 1, we first split the RGB three channels of the original input image to create three single-channel RGB images. Then we apply DWT to each single-channel image, and retain the LL subband images for normalization. The GT images and the preprocessed grayscale images are summed at pixel wise as the corresponding ground truth images. Since the sizes of the retained LL subband images are $112\times 112$ (half of the $224\times 224$ length and width), we resize the images to the designated size of $224\times 224$ using bicubic interpolation [32].
3.2.1 Discrete Wavelet Transform
In this work, we use Haar wavelet transform [33], which is one of the DWT, to process the input image. This process can be conceived as a bandpass filter, selectively allowing signals with frequencies akin to those of the wavelet basis function to traverse. The flowchart of this DWT processing is illustrated in Figure 2, which generates four subband images: one low-frequency subband image and three high-frequency subband images. This work takes the single-channel split images as input and processes the n-direction by low-pass filtering, as shown in the following equation:
$\begin{split}v_{L}[m,n]=\sum_{k=0}^{K-1}x[m,2n-k]g[k]\end{split}$ (2)
Next, this work uses the low-pass filtering applied to the m-direction of Eq. (2) to obtain the following equation:
$\begin{split}v_{LL}[m,n]=\sum_{k=0}^{K-1}v_{L}[2m-k,n]g[k]\end{split}$ (3)
In contrast to the ground truth images provided by the dataset, the RGB single-channel images split from the original input images exhibit some substantial noise. This noise prevents them from being used directly in generating corresponding ground truth images through the generator and could potentially disrupt the performance of the trained adversarial network. DWT decomposes the image into both low-frequency and high-frequency information. The low-frequency component represents the mean value and the high-frequency component represents the difference value. The mean value, classified as low-frequency information, captures gradual changes and encodes both the picture’s contour information and approximate details. Conversely, the difference value, characterized as high-frequency information, captures rapid changes and encodes intricate picture details as well as local information (including noises). Therefore, we apply DWT and retain the LL subband images to input Stage-2, as shown in Figure 2, which effectively preserves contour information and reduces noise interferences.
Figure 3: Results of Haar wavelet transform & normalization of RGB single-channel images split from the example image patches.
Figure 4: Image normalization from image HW15 of DIBCO 2011 dataset, the above is an example of no normalization after DWT retains LL subband images, whereas the below is the example of normalization after DWT retains LL subband images.
3.2.2 Normalization
Normalization entails constraining data within a desired range, with the goal of simplifying the data processing and expediting convergence. The proposed equation for this process is expressed as:
$\begin{split}I_{N}=(\operatorname{IV_{nmax}}-\operatorname{IV_{nmin}})\frac{1}{1+e^{-\frac{I-\beta}{\alpha}}}+\operatorname{IV_{nmin}}\end{split}$ (4)
where IV is the intensity values, the image to be processed $I:\left\{\mathbb{X}\subseteq\mathbb{R}^{n}\right\}\rightarrow\{\operatorname{IV_{min}},\ldots,\operatorname{IV_{max}}\}$, and the generated image $I_{N}:\left\{\mathbb{X}\subseteq\mathbb{R}^{n}\right\}\rightarrow\{\operatorname{IV_{nmin}},\ldots,\operatorname{IV_{nmax}}\}$. $\alpha$ defines the width of the input intensity range, and $\beta$ defines the range centered intensity.
In the proposed method, we normalize the LL subband images to present initially incomparable data for comparison. The reason we choose the normalized LL subband image as the output is shown in Figure 3, where we apply DWT and normalization to each subband images.
The illustrative example of the impact of normalization on RGB single-channel images after the DWT operation is provided in Figure 4, when applied to image HW15 from the DIBCO 2011 dataset, normalization effectively balances the RGB. In contrast, without applying normalization results in an output image, it heavily skews towards the blue component. In addition, we also show that the global binarization results of the images using normalization are more similar to the Ground Truth images provided by the dataset compared to the images only processed by DWT, and the experiments are explained in detail in Section 4.5.
Figure 5: The architecture of Stage-2 of the proposed method. This method employs the U-Net++ [34, 35] architecture with EfficientNet-B6 [36] as the generator and improved PatchGAN [26] as the discriminator. The ground truth image during the training process is obtained from pixel-wise summation of the dataset provided image and the output from Stage-1.
Figure 6: The architecture of Stage-3 of the proposed method. This stage comprises both local and global predictions. The local prediction involves training several $224\times 224$ image patches, while the global prediction involves directly resizing the original image to $512\times 512$ for training.
3.3 Stage-2
In Stage-2, the proposed method is illustrated in Figure 5. Initially, the original input images are split into three RGB single-channel images and one grayscale image. Each of the single-channel images is input into an independent generator, resulting in the use of four generators in Stage-2 of the proposed method. These four generators train the model separately, enabling this method to remove background information from the localized image blocks while extracting colored foreground details. Additionally, all single-channel images share the same discriminator to distinguish between the generated image and its corresponding ground truth image.
3.3.1 Loss functions of GAN
To ensure a more robust convergence of the loss function, we follow the improved Wasserstein generative adversarial network with gradient penalty (WGAN-GP) [37] used by Suh et al. [31] applied to the objective function. Experiments by Bartusiak et al. [38] have demonstrated that the binary cross-entropy (BCE) loss yields superior performance compared to the $L1$ loss in binary classification tasks. Therefore, we chose the BCE loss instead of the $L1$ loss used in the Pix2Pix GAN [26]. The loss function of the WGAN-GP, which incorporates the BCE loss, is formulated as follows:
$\begin{split}\mathop{\mathbb{L}_{D}}=-\mathop{\mathbb{E}_{x,y}}[D(y,x)]+\mathop{\mathbb{E}_{x}}[D(G(x),x)]\\ +\alpha\mathop{\mathbb{E}_{x,\hat{y}\sim P_{\hat{y}}}}[(||\nabla_{\hat{y}}D(\hat{y},x)||_{2}-1)^{2}]\end{split}$ (5)
$\begin{split}\mathop{\mathbb{L}_{G}}=\mathop{\mathbb{E}_{x}}[D(G(x),x)]+\lambda\mathop{\mathbb{E}_{G(x),y}}[y\log G(x)\\ +(1-y)\log(1-G(x))]\end{split}$ (6)
where the penalty coefficient is $\alpha$, and the uniform sampling along a straight line between the ground truth distribution $P_{y}$ and the point pairs of the generated data distribution is $P_{\hat{y}}$. $\lambda$ is used to control the relative importance of different loss terms. In the discriminator, the generated image is distinguished from the real image by the loss function $\mathbb{L}_{D}$ in Eq. (5). In the generator, the distance between the generated image and the ground truth image in each color channel is minimized by the loss function $\mathbb{L}_{G}$ in Eq. (6).
3.3.2 Network architecture
Given the unpredictability of color document degradation, this work adopts four independent adversarial networks to extract text information from diverse colored backgrounds. This method mitigates color interference during document image binarization.
For the network proposed in Stage-2, four independent generators apply the same encoder-decoder architecture. The encoder extracts features, while a shared decoder is employed by multiple encoders of varying depths, and subsequently the features are concatenated. A skip connection is used between the encoder and the decoders. The encoder facilitates downsampling and context extraction, while the decoders undertake upsampling and amalgamate both upscaled and downsampled features. Specifically, this work employs EfficientNet-B6 [36] as the encoder, which is a lightweight CNN known for its performance in image classification tasks. For the discriminator, we choose the one [26, 29, 30, 39, 40, 41] used by Suh et al. [31] due to its superior generalization capabilities. Additionally, instead of U-Net [21] used by Suh et al. [31], this work uses its improved model U-Net++ [34, 35] for semantic segmentation, which can further improve the model performance.
The detailed procedure for training the proposed GAN models is illustrated in Algorithm 1. The training process is conducted in our training environment, which is described in Section 4.3, encompassing optimizer selection, learning rate configuration, and loss function definition. The algorithm demonstrates that the GAN model iteratively updates the loss function based on the authenticity assessment of the generated image by the discriminator. This iterative process enhances the ability of the generator to produce more real images.
Algorithm 1 Training process for the proposed GAN models
0:  Input images $I_{input}$
0:  $\omega=0.5$, $\alpha=10$, $\lambda=50$, $\eta=1\times 10^{-4}$
1:  Initialize: Parameters $\theta_{G_{r}}$, $\theta_{G_{g}}$, $\theta_{G_{b}}$, $\theta_{G_{gray}}$
2:  for number of training iterations do
3:     Split $I_{input}$ into $\{red,green,blue,gray\}$ images
4:     for $k=\{red,green,blue,gray\}$ do
5:        if $k$ is gray then
6:           $y_{k}\leftarrow y$
7:        else
8:           $x^{\prime}_{k}=Norm(V_{LL}(x_{k}))$
9:           $y_{k}\leftarrow x^{\prime}_{k}\bigcap y$
10:           Binarize $y_{k}$ with threshold value $t$.
11:        end if
12:        Update Discriminator $D$ (Eq. (5)):
13:        $\theta_{D}\leftarrow\theta_{D}-\eta_{D}\nabla_{\theta_{D}}\mathop{\mathbb{L}_{D}}$
14:        Update Generator $G_{k}$ (Eq. (6)):
15:        $\theta_{G_{k}}\leftarrow\theta_{G_{k}}-\eta_{G}\nabla_{\theta_{G_{k}}}\mathop{\mathbb{L}_{G}}$
16:     end for
17:  end for
18:  for $k=\{red,green,blue\}$ do
19:     $\hat{y}_{k}\leftarrow\omega G_{k}(x_{k})+(1-\omega)G_{gray}(x_{gray})$
20:  end for
21:  $\hat{y}\leftarrow[\hat{y}_{r},\hat{y}_{g},\hat{y}_{b}]$
3.4 Stage-3
In Stage-3, a multi-scale adversarial network is employed to generate images for both local and global binarization results, enhancing the distinction between the background and text with higher accuracy. Contrary to Stage-2, which predominantly relies on local prediction using the image patches ($224\times 224$), Stage-3 conducts global prediction on the resized input image ($512\times 512$). This method reduces the contextual information loss in text caused by the local predictions of Stage-2.
Figure 6 illustrates the process of two independent discriminators in Stage-3. Different from Stage-2 using 8-bit images for local prediction, Stage-3 uses 24-bit images as the input of the generator. Additionally, Figure 6 clarifies that the left-side input image comprises the image patches from the Stage-2 output, intended for local prediction. Conversely, the right-side input image represents the resized image from the original input for global prediction. In the binarization process of Stage-3, the network architecture of the generator remains consistent with Stage-2, differing only in the number of input channels.
4 Experiments
4.1 Datasets
4.1.1 DIBCO
The Document Image Binarization Competition (DIBCO) provides several datasets: DIBCO 2009 [42], H-DIBCO 2010 [43], DIBCO 2011 [11], H-DIBCO 2012 [44], DIBCO 2013 [12], H-DIBCO 2014 [13], H-DIBCO 2016 [14], DIBCO 2017 [15], and H-DIBCO 2018 [16]. These datasets comprise both grayscale and color images, including machine-printed as well as handwritten content. Participants in the competition were tasked with designing algorithms capable of extracting binarized text images from both handwritten and machine-printed images within these datasets.
4.1.2 PHIBD
The Persian Heritage Image Binarization Dataset (PHIBD) [45] contains 15 images of historical manuscripts from Mirza Mohammad Kazemaini’s old manuscript library in Yazd, Iran. The images within the dataset have suffered from various types of degradation.
4.1.3 SMADI
The Synchromedia Multispectral Ancient Document Images Dataset (SMADI) [1] consists of 240 multispectral images of 30 authentic historical handwritten letters. These letters, dating from the 17th to the 20th centuries, were written using iron gall ink. The original documents are deposited at the BAnQ Bibliotheque et Archives nationales du Quebec. A total of 240 images of real documents were meticulously captured, calibrated, and registered within this dataset. Each of these images was imaged with a CROMA CX MSI camera, resulting in 8 images per document.
4.1.4 Bickley Diary Dataset
The Bickley Diary Dataset (BD) [46] has been generously donated to the Singapore Methodist Archives by Erin Bickley. This dataset comprises diaries in which the images have been affected by both light and fold damages, thereby complicating the process of identification.
4.2 Evaluation Metric
In comparison with other traditional and SOTA methods for document image binarization, based on the case that DIBCO datasets do not provide OCR output, we use the following five evaluation metrics:
4.2.1 F-measure (FM)
${\rm FM}=\frac{2\times{\rm Recall}\times{\rm Precision}}{{\rm Recall}+{\rm Precision}}$ (7)
where the Recall is defined as $\frac{TP}{TP+FN}$; the Precision is defined as $\frac{TP}{TP+FP}$, and TP, FP, and FN represent the true positive, false positive, and false negative values, respectively.
4.2.2 Pseudo-F-measure (p-FM)
${\rm p-FM}=\frac{2\times{\rm Recall_{ps}}\times{\rm Precision_{ps}}}{{\rm Recall_{ps}}+{\rm Precision_{ps}}}$ (8)
where the Pseudo-Recall (${\rm Recall_{ps}}$) and the Pseudo-Precision (${\rm Precision_{ps}}$) both use the distance weights relative to the ground truth image profile to generate a new weighted ground truth image.
4.2.3 Peak signal-to-noise ratio (PSNR)
${\rm PSNR}=10\log(\frac{{\rm V}^{2}}{{\rm MSE}})$ (9)
where ${\rm V}$ represents the difference between the foreground and background, and the Mean Square Error (MSE) is defined as:
${\rm MSE}=\frac{\sum_{x=0}^{M-1}\sum_{y=0}^{N-1}{[}L(x,y)-L^{\prime}(x,y){]}^{2}}{M\times N}$ (10)
where $M$ and $N$ are the total number of pixels of image $L$ and image $L^{\prime}$ respectively. Generally, a higher PSNR value indicates greater similarity between the images.
4.2.4 Distance reciprocal distortion (DRD)
${\rm DRD}$ represents a measure of visual distortion in the binary image, which is shown in the following equation:
${\rm DRD}=\frac{{\rm\sum_{k}DRD_{k}}}{{\rm NUBN}}$ (11)
where ${\rm NUBN}$ corresponds to the number of uneven $8\times 8$ blocks in the ground truth image, and ${\rm DRD_{k}}$ is defined as:
${\rm DRD_{k}}=\sum_{i=-2}^{2}\sum_{j=-2}^{2}|{\rm G_{k}(i,j)-B_{k}(x,y)|\times N_{w}(i,j)}$ (12)
where ${\rm DRD_{k}}$ indicates the distortion of the ${\rm k_{th}}$ flipped pixel, calculated using a $5\times 5$ normalized weight matrix ${\rm N_{w}}$. ${\rm B_{k}}$ refers to the pixel of the binary image, and ${\rm G_{k}}$ denotes the pixel of the ground truth image.
4.2.5 Avg-Score (Avg)
The Peak Signal-to-Noise Ratio (PSNR) provides a straightforward and intuitive quantitative metric for comparing the generated image with the ground truth image. Although it can be used as a supplementary metric, it should not be the only evaluation measure. The F-measure and p-F-measure are used to assess the accuracy of different methods, but a higher value in these metrics does not always correspond to a higher PSNR value. In our experiments, it is common to observe that a method achieves the SOTA PSNR value, while the F-measure and p-F-measure values are lower than those of other models. To solve this problem, we followed the Avg-Score metric used by Jemni et al. [47] to compare the combined performance of different methods, as shown in the following equation:
$Avg=\frac{FM+p\!-\!FM+PSNR+(100-DRD)}{4}$ (13)
Table 1: Comparison of the PSNR values of the images from different training sets that have been processed using different methods with the corresponding ground truth images.
Method DIBCO 2009 H-DIBCO 2010 H-DIBCO 2012 Bickley Diary PHIBD SMADI Mean Values
Original Input 63.2879 61.3446 61.8542 57.6995 60.7355 61.2175 61.0232
DWT 64.1154 71.4422 64.8151 53.7037 51.5094 60.1374 60.9539
DWT&Norm 80.4286 86.1904 82.4625 74.2314 83.6403 78.5337 80.9145
Table 2: Quantitative comparison (FM/p-FM/PSNR/DRD/Avg) of the models trained with different patch sizes ($224\times 224$/$256\times 256$) on multiple benchmark datasets.
DIBCO 2011 DIBCO 2013 H-DIBCO 2014 H-DIBCO 2016 DIBCO 2017 DIBCO 2018 Mean Values
$256\times 256$ FM↑ 94.11 95.06 96.62 91.40 90.95 91.76 93.32
p-FM↑ 96.97 97.30 98.01 96.22 93.97 95.74 96.37
PSNR↑ 20.48 22.16 22.25 19.63 18.55 19.98 20.51
DRD↓ 1.76 1.66 0.96 2.96 2.95 2.75 2.17
Avg↑ 77.45 78.22 78.98 76.07 75.13 76.18 77.01
$224\times 224$ FM↑ 94.22 94.65 96.60 91.76 91.31 92.89 93.57
p-FM↑ 97.47 97.10 98.33 96.82 94.30 96.96 96.83
PSNR↑ 20.54 22.01 22.23 19.73 18.63 20.39 20.59
DRD↓ 1.69 1.98 0.97 2.81 2.88 2.23 2.09
Avg↑ 77.64 77.95 79.05 76.38 75.34 77.00 77.22
Figure 7: Visualization examples of the results of the proposed method at each stage. (a) the original input images, (b) the LL subband images obtained by Haar wavelet transform and normalization (Stage-1), (c) the enhanced images produced using our image enhancement method (Stage-2), (d) the binarized images obtained by combining local and global predictions (Stage-3), (e) the ground truth images.
Table 3: Comparison of performance evaluation of the proposed method and other methods for document binarization on multiple benchmark datasets. Best and 2nd best performance are in red
and blue colors, respectively. The data for the experimental results of other methods are obtained from [31, 48].
Method FM↑ p-FM↑ PSNR↑ DRD↓ Avg↑
Niblack [5] 70.44 73.03 12.39 24.95 57.73
Otsu [6] 78.60 80.53 15.31 22.57 62.97
Sauvola [7] 85.37 89.08 16.37 7.08 70.94
Howe [17] 94.04 95.06 20.43 2.10 76.86
Jia [18] 93.17 95.32 19.44 2.29 76.41
Vo [23] 92.58 94.67 19.16 2.38 76.01
He [24] 91.92 95.82 19.49 2.37 76.22
Suh [31] 93.57 95.93 20.22 1.99 76.93
Ours 94.22 97.47 20.54 1.69 77.64
(a) DIBCO 2011
Method FM↑ p-FM↑ PSNR↑ DRD↓ Avg↑
Niblack [5] 71.38 73.17 13.54 23.10 58.75
Otsu [6] 80.04 83.43 16.63 10.98 67.28
Sauvola [7] 82.71 87.74 17.02 7.64 69.96
Howe [17] 91.34 91.79 21.29 3.18 75.31
Jia[18] 93.28 96.58 20.76 2.01 77.15
Vo [23] 93.43 95.34 20.82 2.26 76.83
He [24] 93.36 96.70 20.88 2.15 77.20
Suh [31] 95.01 96.49 21.99 1.76 77.93
Ours 94.65 97.10 22.01 1.98 77.95
(b) DIBCO 2013
Method FM↑ p-FM↑ PSNR↑ DRD↓ Avg↑
Niblack [5] 86.01 88.04 16.54 8.26 70.58
Otsu [6] 91.62 95.69 18.72 2.65 75.85
Sauvola [7] 84.70 87.88 17.81 4.77 71.41
Howe [17] 96.49 97.38 22.24 1.08 78.76
Jia [18] 94.89 97.68 20.53 1.50 77.90
Vo [23] 95.97 97.42 21.49 1.09 78.45
He [24] 95.95 98.76 21.60 1.12 78.80
Suh [31] 96.36 97.87 21.96 1.07 78.78
Ours 96.60 98.33 22.23 0.97 79.05
(c) H-DIBCO 2014
Method FM↑ p-FM↑ PSNR↑ DRD↓ Avg↑
Niblack [5] 72.57 73.51 13.26 24.65 58.67
Otsu [6] 86.59 89.92 17.79 5.58 72.18
Sauvola [7] 84.64 88.39 17.09 6.27 70.96
Howe [17] 87.47 92.28 18.05 5.35 73.11
Jia [18] 90.10 93.72 19.00 4.03 74.70
Vo [23] 90.01 93.44 18.74 3.91 74.57
He [24] 91.19 95.74 19.51 3.02 75.86
Suh [31] 92.24 95.95 19.93 2.77 76.34
Ours 91.76 96.82 19.73 2.81 76.38
(d) H-DIBCO 2016
Method FM↑ p-FM↑ PSNR↑ DRD↓ Avg↑
Otsu [6] 77.73 77.89 13.85 15.54 63.48
Sauvola [7] 77.11 84.10 14.25 8.85 66.65
Howe [17] 90.10 90.95 18.52 5.12 73.61
Jia [18] 85.66 88.30 16.40 7.67 70.67
Suh [31] 90.95 94.65 18.40 2.93 75.27
Ours 91.31 94.30 18.63 2.88 75.34
(e) DIBCO 2017
Method FM↑ p-FM↑ PSNR↑ DRD↓ Avg↑
Otsu [6] 51.45 53.05 9.74 59.07 38.79
Sauvola [7] 67.81 74.08 13.78 17.69 59.50
Howe [17] 80.84 82.85 16.67 11.96 67.10
Jia [18] 76.05 80.36 16.90 8.13 66.30
Suh [31] 91.86 96.25 20.03 2.60 76.39
Ours 92.89 96.96 20.39 2.23 77.00
(f) DIBCO 2018
Figure 8: Example of document image binarization results from input image HW1 of DIBCO 2011 dataset by different methods: (a) original input, (b) Otsu [6], (c) Niblack [5], (d) Sauvola [7], (e) Vo [23], (f) He [24], (g) Suh [31], (h) Ours, (i) ground truth.
Figure 9: Example of document image binarization results from input image HW5 of DIBCO 2013 dataset by different methods: (a) original input, (b) Otsu [6], (c) Niblack [5], (d) Sauvola [7], (e) Vo [23], (f) He [24], (g) Suh [31], (h) Ours, (i) ground truth.
4.3 Implementation Details
4.3.1 Data preparation
The performance of a neural network depends on the quality and quantity of its training data. In general, the more training data enhances the better the performance of the neural network. To compare the performance of the proposed method with other methods, we follow the strategy described in [23, 24, 31] for constructing the training and test sets. Specifically, the training set includes a total of 143 images from DIBCO 2009 [42], H-DIBCO 2010 [43], H-DIBCO 2012 [44], PHIBD [45], SMADI [1], and BD [46] dataset. The test sets comprise a total of 82 images from DIBCO 2011 [11], DIBCO 2013 [12], H-DIBCO 2014 [13], H-DIBCO 2016 [14], DIBCO 2017 [15], and H-DIBCO 2018 [16].
To ensure a fair comparison of performance among all methods, we used the same data augmentation for our method and other SOTA methods. During the training process of the local prediction network, this work splits the input images into patches with $224\times 224$ in size. These patches were sampled with scale factors of ${0.75,1.25,1.5}$, subsequently resized to $224\times 224$, and rotated by 270°. In contrast, for training the global prediction network, this work employed scaling without rotation for data augmentation, using a scaling size of $512\times 512$. To diversify the training data, rotation augmentation was applied to the global prediction, involving rotations of 90°, 180°, and 270°, as well as horizontal and vertical flips.
4.3.2 Training
The experiments were conducted using the PyTorch framework in Python. We use NVIDIA RTX 3090 GPUs for all the following experiments.
In this study, we employ the generator comprising EfficientNet-B6 [36] with UNet++ [34, 35]. Regarding the training hyperparameters, Stage-2 and Stage-3 share similar settings, differing primarily in the number of training epochs: 150 epochs for global prediction and 10 epochs for the remaining networks. The choice of optimizer was Adam [49], with a learning rate set at $2\times 10^{-4}$. Additionally the generator employed $\beta_{1}=0.5$, while the discriminator used $\beta_{2}=0.999$.
4.3.3 Pre-training
To enhance training efficacy due to limited data availability, we employed EfficientNet-B6 [36] as the encoder in the generator, utilizing weights pretrained on the ImageNet dataset [50].
4.4 Comparison with other Methods
In our previous conference paper [51], we compared various deep learning-based methods and found that some of these methods [10, 28, 29] employed a “leave-one-out” strategy for selecting the training set. For instance, when the DIBCO 2013 dataset is used as the test set, all other datasets are utilized as the training set. Although our method still demonstrates superior performance at the SOTA level in such case [51], we consider it unfair to compare them directly because our model uses a smaller training set. Therefore, this work excludes these methods from this paper and only retains those that use the exact same training set for comparison.
4.5 Image Enhancement Experiment
In Stage-1, we applied DWT and normalization to the original input images. While mathematical theory [52] suggests that these processed images can retain contour information and reduce noise more efficiently, it is important to demonstrate their impact on experimental results. The previous conference paper [51] on this work presented ablation experiments on DWT and normalization. Based on this, it is better to intuitively show that, compared to the original input images, the DWT and normalization processed images are closer to the ground truth images.
In Stage-1 of this work, we set the threshold to 0.3 to binarize different single-channel images to obtain corresponding binarized images. In the experiment of this section, we applied the above image binarization to the original input images, the images processed with DWT, and the images processed with both DWT and normalization. Subsequently, we calculated the PSNR values of the binarized images compared to the ground truth images using these three methods, and the results are presented in Table 2. The PSNR mean values of the original input images are 61.0232. In contrast, the PSNR mean values of the images after DWT processing are slightly lower at 60.9539. However, the PSNR mean values of the images processed with both DWT and normalization are 80.9145, which means they are closer to the ground truth images. This demonstrates that it is more reasonable to choose the DWT and normalization processed image rather than the original input image and the DWT processed image.
4.6 Comparison Experiment
To ensure fairness in the experiments, we used the same training set and data augmentation techniques described in Section 4.3.1 for all the compared models. Unlike Suh et al. [31], who split the input images into $256\times 256$ patches, our proposed method uses $224\times 224$ patches. To demonstrate the positive impact of this change on the model performance, we conducted experiments comparing models using $224\times 224$ patches with those using $256\times 256$ patches, keeping all other settings unchanged.
Table 2 presents the performance of the models across different patch sizes on six DIBCO datasets. As established in Section 4.5, the images processed with DWT and normalization outperform those processed with either the original input images or only with DWT. Consequently, all models in this section utilize images processed with DWT and normalization. According to Table 2, the mean values of FM, p-FM, PSNR, DRD, and Avg for the model based on $224\times 224$ patches are 93.57, 96.83, 20.59, 2.09, and 77.22, respectively, which are superior to those for the model based on $256\times 256$ patches. This indicates that the model based on $224\times 224$ patches performs better, and therefore we will use this model for subsequent comparisons.
4.7 Experimental Results
Different from the conference paper [51] of this work, which uses option 3 to train the model, the proposed method in this paper applies DWT and normalization to the images in Stage-1 as input to train the model. The performance of the trained model will then be compared with that of other methods.
Figure 7 presents the visualization results at each stage of the trained model. Specifically, five images were randomly selected from the PHIBD and Bickley Diary datasets to demonstrate the step-by-step process of image enhancement and binarization using the proposed method. Figure 7 (b) represents the result of retaining the LL subband images after applying DWT and normalization. This stage performs noise reduction on the original input image, which is used as the ground truth image for the generator. Figure 7 (c) shows the result of image enhancement achieved using adversarial networks, where the image has the background color removed and the color of the text highlighted. Figure 7 (d) is the final output image obtained using the proposed whole method. Notably, it is evident that our final output result closely approximates Figure 7 (e), the ground truth image.
Table 3 presents the evaluation results of different models on six DIBCO datasets. Overall, the proposed method achieves the top two best performance in terms of FM, p-FM, PSNR, and DRD, with its Avg value reaching the SOTA level across all six datasets. Notably, the performance of the proposed method on the DIBCO 2011 and 2013 datasets is particularly good, obtaining the best values for each of the evaluated metrics. For instance, on the DIBCO 2011 dataset, the p-FM value of the proposed method is 97.47, significantly surpassing the second best value of 95.93. Similarly, on the DIBCO 2018 dataset, the FM value of the proposed method is 92.89, considerably higher than the second best value of 91.86. These results indicate that the proposed method is more effective than other methods for the binarization of degraded color documents.
In addition to assessing method performance by comparing quantitative metrics across various methods, this work aims to provide readers with an intuitive understanding of the distinctions by the visualization results. Consequently, we have chosen images from the DIBCO 2011 and 2013 datasets to illustrate examples of binarization results. Figures 8 and 9 show the binarization results of images HW1 and HW5 by different methods. These figures visually demonstrate that the proposed method outperforms deep learning based methods and the traditional binarization methods in terms of the shadow and noise elimination. Furthermore, the proposed method is good at preserving textual content while effectively mitigating the presence of the shadows and noise.
5 Conclusion
This paper introduces a novel three-stage method based on DWT and GANs to apply the task of text degradation in degraded document images with color backgrounds. The proposed method involves preserving the LL subband images through DWT to eliminate image noise, serving as the ground truth images for the generator. It trains independent networks utilizing RGB three single-channel images and one grayscale image, followed by a fusion of local and global prediction networks for binarization. As far as our knowledge extends, this work first represents the three-stage method that combines traditional image processing with deep learning for document binarization. The experimental results demonstrate that employing U-Net++ with EfficientNet-B6 as the generator and PatchGAN as the discriminator outperforms performance compared to other SOTA methods.
6 Declarations
This paper is an expanded paper from 20th Pacific Rim International Conference on Artificial Intelligence (PRICAI) held on November 15-19, 2023 in Jakarta, Indonesia.
6.1 Declaration of Funding
This research is supported by National Science and Technology Council of Taiwan, under Grant Number: NSTC 112-2221-E-032-037-MY2.
6.2 Declaration of competing interest
The authors have no financial or proprietary interests in any material discussed in this article.
6.3 Declaration of Generative AI and AI-assisted technologies in the writing process
The authors only use generative artificial intelligence (AI) and AI-assisted technologies to improve language.

### v8

@@
 Abstract
-The efficient segmentation of text information from the background in degraded color document images is an important challenge in the preservation of ancient manuscripts.
-The imperfect preservation of ancient manuscripts has led to various types of degradation over time, such as staining, yellowing, and ink seepage, badly affecting document image binarization results.
-This work proposes a three-stage method to generate binarization image results for degraded colour document images using generative adversarial networks (GANs).
-Stage-1 involves applying discrete wavelet transform and retaining the low-low subband images for document image enhancement.
-In Stage-2, the original input image is split into red, green, and blue (RGB) three single-channel images and one grayscale image, and each image is trained with independent GANs to extract color foreground information.
-In Stage-3, the output images of Stage-2 and the resized input images are used to train independent GANs to generate document binarization results, enabling the combination of global and local features.
-The experimental results show that the Avg-Score of the proposed method is 77.64, 77.95, 79.05, 76.38, 75.34, and 77.00 on the (H)-DIBCO 2011, 2013, 2014, 2016, 2017, and 2018 datasets, which achieves the state-of-the-art level.
+The efficient extraction of text information from the background in degraded color document images is an important challenge in the preservation of ancient manuscripts.
+The imperfect preservation of ancient manuscripts has led to different types of degradation over time, such as page yellowing, staining, and ink bleeding, seriously affecting the results of document image binarization.
+This work proposes an effective three-stage network method to image enhancement and binarization of degraded documents using generative adversarial networks (GANs).
+Specifically, in Stage-1, we first split the input images into multiple patches, and then split these patches into four single-channel patch images (gray, red, green, and blue).
+Then, three single-channel patch images (red, green, and blue) are processed by the discrete wavelet transform (DWT) with normalization.
+In Stage-2, we use four independent generators to separately train GAN models based on the four channels on the processed patch images to extract color foreground information.
+Finally, in Stage-3, we train two independent GAN models on the outputs of Stage-2 and the resized original input images ($512\times 512$) as the local and global predictions to obtain the final outputs.
+The experimental results show that the Avg-Score metrics of the proposed method are 77.64, 77.95, 79.05, 76.38, 75.34, and 77.00 on the (H)-DIBCO 2011, 2013, 2014, 2016, 2017, and 2018 datasets, which are at the state-of-the-art level.
 The implementation code for this work is available at https://github.com/abcpp12383/ThreeStageBinarization.
@@
 1 Introduction
-Historical analyses of human societies primarily rely on reconstructing and analyzing preserved ancient documents.
-Parchment manuscripts and paintings have predominantly served as the means of preserving crucial ancient documents throughout the evolution of human society.
-Analyzing preserved ancient documents presents challenges due to potential forms of interference and degradation, including paper yellowing, text fading, page contamination and moisture damage.
-Document image enhancement is one of the most important tasks in the reconstruction and analysis of ancient documents.
-It involves document image binarization to extract essential text information from the page.
-Elevating the image quality of deteriorated documents necessitates addressing various challenges, including text degradation and bleed-through issues, such as text degradation and bleed-through [1, 2, 3, 4].
-However, traditional binarization methods cannot extract text information from the background of degraded documents effectively, often resulting in incomplete removal of shadows and noises, or loss of text information [5, 6, 7].
+Historical analysis of ancient human society mainly relies on the reconstruction and analysis of preserved ancient documents, where the parchment manuscripts are an important tool for preserving important ancient documents.
+Analyzing preserved ancient documents, especially the parchment manuscripts, presents significant challenges due to potential forms of degradation, including paper yellowing, text fading, page contamination, and moisture damage [1, 2].
+Document image enhancement and binarization are important steps in analyzing ancient documents, with the purpose of extracting useful text information from the background.
+However, improving the image quality of degraded documents requires solving several problems, such as text degradation and ink bleeding [3, 4].
+Traditional threshold methods cannot effectively extract text information from the background of degraded documents, often resulting in incomplete removal of shadows and noise, or loss of text information [5, 6, 7].
 With the development of deep learning, neural networks have been applied to image binarization.
-Calvo-Zaragoza and Gallego [8] introduced a selectional auto-encoder approach, parsing the text through the trained network model, and performing binarization using threshold.
-Fully convolutional network (FCN) [9] enables the generated images to maintain the same size as the input images, and the task [10] of generating document binarization using FCN is also included in the semantic segmentation task.
-However, most image binarization methods primarily process grayscale documents, targeting contaminated black and white scanned ancient documents as their default.
-Considering that the images of some ancient documents are in colour, we propose a novel three-stage method for document image enhancement and colour degradation document binarization.
-The proposed method consists of three stages.
-In Stage-1, we split the original input image into red, green, and blue (RGB) three single-channel images and one grayscale image, and apply discrete wavelet transform (DWT) to retain the low-low (LL) subband images and normalize them to achieve image enhancement.
-In Stage-2, we train four independent adversarial networks with the four single-channel images to eliminate background information and extract color foreground details from the local image patches.
-During the processing of Stage-3, the multiscale adversarial networks produce $224\times 224$ image patch (local) prediction outputs and $512\times 512$ resized image (global) prediction outputs for the document images, which are subsequently fused to obtain the final output.
-To evaluate the performance of the proposed method for degraded document image binarization, we compare it with other methods using DIBCO datasets of 2011, 2013, 2014, 2016, 2017, and 2018 [11, 12, 13, 14, 15, 16].
+The common method is to apply neural networks to analyze the document images, for instance, Calvo-Zaragoza and Gallego [8] introduce a selectional auto-encoder approach to analyze document images through the trained network model, and binarize the document images using the threshold.
+The classic network in semantic segmentation tasks, fully convolutional network (FCN) [9], can also be used to segment foreground and background information in document images.
+Tensmeyer et al.’s [10] research also shows that using FCN for document binarization can achieve good results.
+However, most current methods mainly process gray-scale document images, targeting contaminated black and white scanned ancient documents as their default.
+Considering some ancient document images are in color, we propose a novel three-stage network method for both gray-scale and color degraded document image enhancement and binarization.
+Specifically, the proposed method consists of three stages.
+In Stage-1, we first split the input images into multiple patches, and then split these patches into four single-channel patch images (gray, red, green, and blue).
+After that, we apply the discrete wavelet transform (DWT) on three single-channel patch images (red, green, and blue) to retain the low-low (LL) subband images, and then normalize them.
+To remove the noise from the background of the original input images, and to extract the foreground information from the patch images, in Stage-2, we train the GANs model on four single-channel images by using four independent generators to generate the enhanced images.
+In addition, in Stage-3, we use the entire document images composed of the patch images (output of Stage-2) as the local predictions, and the $512\times 512$ resized input images as the global predictions to train two independent GAN models, respectively.
+The outputs of these two GAN models are fused to obtain the final output images.
+To evaluate the performance of the proposed method for document image binarization, we compare it with other methods using the (H)-DIBCO datasets of 2011, 2013, 2014, 2016, 2017, and 2018 [11, 12, 13, 14, 15, 16].
 The main contributions of this paper are as follows:
-  • Introducing a novel three-stage generative adversarial networks method for enhancing and binarizing the degraded color documents.
-  • The incorporation of DWT prior to the training of generative adversarial networks makes the image to be input closer to the Ground Truth image.
-  • The proposed method achieves the state-of-the-art performance on multiple benchmark datasets.
-The rest of this paper is organized as follows: Section 2 describes the deep learning method for document image binarization.
-Section 3 introduces a novel three-stage generative adversarial networks method.
-Section 4 presents the experiments performed on this work, and compares the performance of our method and other methods on several benchmark datasets.
-Finally, Section 5 discusses the conclusion and the future work of this paper.
+  • Introduces an effective three-stage network method that effectively performs enhancement and binarization on gray-scale and color degraded document images.
+  • Shows that, compared with other methods using the same training set, the proposed method achieves the state-of-the-art performance on multiple benchmark datasets.
+  • Demonstrates that for document images, the binarization results of the images processed by the discrete wavelet transform with normalization are closer to the Ground-Truth images compared to the binarization results of the original input images.
+The rest of this paper is organized as follows: Section 2 describes the deep learning methods for document image enhancement and binarization.
+Section 3 details the proposed three-stage network method.
+Section 4 presents the results of the image enhancement experiment and the comparison experiment, and shows the comparative results of the performance of the proposed method and other methods on multiple benchmark datasets.
+Finally, Section 5 discusses the conclusion and future work of this work.
 2 Related Work
-In previous researches focusing on document image binarization, the field is broadly divided into two categories: traditional image processing methods and deep learning-based semantic segmentation methods.
-However, traditional image processing binarization methods are always based on the pixel-level local threshold values, such as global binarization [6] and local adaptive binarization [5, 7].
-These methods are not satisfactory when dealing with documents affected by excessive interference or degradation [17, 18].
-With the advancement of deep learning, the convolutional neural network (CNN) framework [19, 20] has become the predominant backbone of computer vision tasks.
-Within computer vision, semantic segmentation is one of the most important research topics, encompassing document image binarization.
+Document image binarization methods can be broadly divided into two categories: traditional threshold methods and deep learning methods.
+Among them, traditional threshold methods are mainly based on calculating the threshold value at the pixel-level of the image, such as global binarization [6] and local adaptive binarization [5, 7].
+However, these methods have unsatisfactory results for documents affected by excessive interference or degradation [17, 18].
+Neural networks [19, 20] continue to achieve excellent results in computer vision tasks.
+Among them, neural networks used in semantic segmentation tasks are also applied to document image binarization.
 For instance, Tensmeyer et al.
-[10] employed pseudo-F-measure (p-FM) and F-measure (FM) losses during the training of FCN to apply the model to the document image binarization task.
-Ronneberger et al.
-[21] extended FCN, presenting the U-Net architecture, which captures contextual and spatial information through its distinctive U-shaped design.
-Inspired by this, Peng et al.
-[22] introduced an encoder-decoder architecture for document image binarization, where the decoder transforms the low-resolution representation to the original dimensions, resulting in the binarized image.
+[10] apply FCN [9] to document image binarization, and combine pseudo-F-measure (p-FM) and F-measure (FM) losses to train the model.
+Inspired by U-Net [21], Peng et al.
+[22] introduce an encoder-decoder architecture for document image binarization, where the decoder transforms the low-resolution representation to the original dimensions, resulting in the binarized images.
 Vo et al.
-[23] first applied hierarchical deep supervised network (DSN) to document image binarization and achieved the state-of-the-art (SOTA) performance on several benchmark datasets.
+[23] apply hierarchical deep supervised network (DSN) to document image binarization, which achieves the state-of-the-art (SOTA) performance on multiple benchmark datasets.
 He et al.
-[24] proposed an iterative deep learning framework applied to document image binarization and the trained model achieved excellent performance.
-In recent years, generative adversarial networks (GANs) [25] have achieved exciting results in the generation of real-world images.
-Unlike GANs which generate output images from random noise, Isola et al.
-[26] proposed Pix2Pix GAN based on conditional GAN (cGAN) [27] for image-to-image translation.
-To ensure that the generated images are more realistic, Pix2Pix GAN employs generators for loss function optimization.
-Within the network architecture, the generator evaluates the minimized generated image against the target output image, quantifying the $L1$ loss between them.
-Meanwhile, the discriminator assesses the credibility of the generated image, relying on the input image and the reference target image.
-This process is depicted by the subsequent equation:
-$\begin{split}\emph{arg}\min_{G}\max_{D}V(G,D)=\mathop{\mathbb{E}_{x}}[\log(1-D(x,G(x)))]\\ +\mathop{\mathbb{E}_{x,y}}[\log D(x,y)]+\lambda\mathop{\mathbb{E}_{x,y}}[||y-G(x)||_{1}]\end{split}$ (1)
-where the input data distribution $\mathop{\mathbb{P}_{x}}$ samples the input image $x$; the ground truth image of the input image is $y$, and the hyperparameter to increase the model regularization effect is $\lambda$.
-Generator $G$ generates the input image as the generated image $G(x)$.
+[24] propose an iterative deep learning framework applied to document image binarization, which also achieves the excellent performance.
+In recent years, GANs [25] have achieved excellent results in image generation tasks.
+Different from generating output images from random noise, Isola et al.
+[26] propose Pix2Pix GAN based on conditional generative adversarial network (cGAN) [27] for image-to-image translation.
+To ensure that the generated images are more realistic, Pix2Pix GAN employs generators to optimize the loss function.
+In the network architecture, the generator evaluates the minimization of generated images against the target output images, quantifying the $L1$ loss between them.
+Meanwhile, the discriminator evaluates the credibility of the generated images based on the input images and the reference target images (the Ground-Truth images).
+The process is described by the following equation:
+$\begin{split}\emph{arg}\min_{G}\max_{D}V(G,D)=\mathop{\mathbb{E}_{x}}[\log(1-D(x,G(x)))]\\ +\mathop{\mathbb{E}_{x,y}}[\log D(x,y)]+\lambda\mathop{\mathbb{E}_{x,y}}[||y-G(x)||_{1}],\end{split}$ (1)
+where the input data distribution $\mathop{\mathbb{P}_{x}}$ samples the input images $x$, and the Ground-Truth images are $y$.
+The hyperparameter to increase the model regularization effect is $\lambda$.
+Generator ($G$) generates the input images as the generated images $G(x)$.
 From Eq.
-(1), the loss function between the generated image and the ground truth image is $L1$ instead of $L2$, which is beneficial to reduce blurring and ambiguity in the generation process.
+(1), the loss function between the generated images and the Ground-Truth images is $L1$ instead of $L2$, which is beneficial to reduce blurring and ambiguity in the generation process.
+Therefore, in addition to neural networks applied in semantic segmentation tasks, GANs have also achieved satisfactory results for document image binarization.
 Bhunia et al.
-[28] performed document image enhancement on the DIBCO dataset and used cGAN [27] for image binarization.
+[28] employ cGAN [27] on the DIBCO datasets for document image enhancement and binarization.
 Inspired by Pix2Pix GAN [26], Zhao et al.
-[29] proposed a cascade generator structure for document image binarization based on Pix2Pix GAN to solve the problem of multi-scale information combination.
+[29] propose a cascade generator structure for document image binarization based on Pix2Pix GAN to solve the problem of multi-scale information combination.
 De et al.
-[30] designed two discriminators to combine local and global information.
+[30] design two discriminators to combine local and global information.
 Suh et al.
-[31] set up a two-stage generative adversarial network for image binarization using Patch GAN [26].
-These methods keep breaking the SOTA performance on DIBCO datasets.
+[31] propose a novel two-stage generative adversarial network method using PatchGANs [26] for document image enhancement and binarization.
+Souibgui et al.
+[32] introduce an efficient end-to-end framework named document enhancement generative adversarial networks (DE-GAN).
+This framework is designed to recover severely degraded document images by generating high-quality enhanced versions of these documents.
+These methods keep breaking the SOTA performance on the DIBCO datasets.
+In addition, diffusion models have also begun to be applied to document image binarization.
+For instance, Cicchetti and Comminiello [33] propose a novel generative framework based on the diffusion probabilistic model (DPM) for document image enhancement, which includes an efficient nonlinear activation free (NAF) network and an additional differentiable module based on the convolutional recurrent neural network (CRNN).
+Figure 1: The overall three-stage network architecture of the proposed method for document image enhancement and binarization.
 Figure 2: The architecture of Stage-1 of the proposed method.
-The original input image is divided into patches of $224\times 224$ size, and split into grayscale images and RGB three single-channel images.
-The single-channel images are all in grayscale.
-However, to visually represent the different channels of the image separately, the single-channel images in this diagram are replaced with images in three different RGB colors.
-Figure 2: The architecture of the Haar wavelet transform used by the proposed method, which takes the red single-channel image patch as an example.
-The processes of the transform of the green and blue single-channel image patches are the same.
+The original input images are split into multiple patches ($224\times 224$), and split into four single-channel patch images (gray, red, green, and blue), where three single-channel patch images (red, green, and blue) are processed by the Haar wavelet transform with normalization.
+Figure 3: The flow chart of the Haar wavelet transform with normalization used by the proposed method, which takes the red single-channel patch images as the example.
+The processes of the green and blue single-channel patch images are the same.
+Figure 4: Example of the Haar wavelet transform with normalization of three single-channel patch images (red, green, and blue).
+Figure 5: Example of the ablation study on image normalization, the above figure presents the image only processed by the Haar wavelet transform, and the below figure presents the image processed by the the Haar wavelet transform with normalization.
+Figure 6: The architecture of Stage-2 of the proposed method.
+This method employs the U-Net++ [34, 35] architecture with EfficientNet-B6 [36] as the generator, and the improved PatchGAN [26] as the discriminator.
+The Ground-Truth images for three single-channel patch images (red, green, and blue) are the binarization results of the outputs from Stage-1.
+Figure 7: The architecture of Stage-3 of the proposed method. This stage comprises the local prediction (left-side) and the global prediction (right-side).
 3 Proposed Method
 3.1 Overall Network Architecture
-Considering the complexity of color document image degradation and the advantages [31] of processing different color channels separately, our proposed method divides the overall network architecture into three stages, which focus on enhancing image quality of degraded color documents and extracting text information from the images.
-In Stage-1, we use DWT and normalization to preprocess the patches of the input image.
-These preprocessed image patches are employed to train as the input of the generator within the adversarial network framework.
-Stage-2 focuses on the elimination of background color noise and the extraction of color foreground information, while Stage-3 involves the fusion of local and global predictions to produce the final output.
-The detailed introduction of each stage can be found in Sections 3.2, 3.3, and 3.4, respectively.
+Considering the complexity of degradation of color documents, better results are often obtained by processing images of different color channels separately [31].
+Therefore, this method divides the overall network architecture into three stages, which are used for document image enhancement, text information extraction, and document image binarization, as shown in Figure 1.
+The images processed in Stage-1 are prepared for training GANs models in Stage-2 and Stage-3.
+Stage-2 focuses on the removal of noise from the background and the extraction color foreground information, while Stage-3 involves the fusion of the local and global predictions to obtain the final output images.
+The detailed introduction of each stage can be found in Sections 3.23.3, and 3.4, respectively.
 3.2 Stage-1
-This work employs four different generators to extract foreground color information from degraded document images while eliminating the background color.
-To preprocess the input of the different generators, we first split the original color image (24-bit) into RGB three single-channel images (8-bit) and one grayscale image (8-bit).
-The ground truth (GT) image provided by the dataset is a single-channel binary image (8-bit), where pixel points have values of 0 for characters and 1 for the background.
-Images with varying channel numbers cannot be directly put into the discriminator for processing.
-As shown in Figure 1, we first split the RGB three channels of the original input image to create three single-channel RGB images.
-Then we apply DWT to each single-channel image, and retain the LL subband images for normalization.
-The GT images and the preprocessed grayscale images are summed at pixel wise as the corresponding ground truth images.
-Since the sizes of the retained LL subband images are $112\times 112$ (half of the $224\times 224$ length and width), we resize the images to the designated size of $224\times 224$ using bicubic interpolation [32].
+This work employs four independent generators to train GANs model, which extracts foreground information from degraded document images while removing noise from the background.
+For color degraded documents, to process the inputs of different independent generators, we first split the original input images (24-bit) into four single-channel patch images (gray, red, green, and blue), each of which is 8-bit.
+However, for the gray-scale (8-bit) degraded documents, we only process the gray single-channel images in Stage-1 and train the model on gray single-channel images for Stage-2.
+Next, we will focus on the case of color degraded documents.
+Since the Ground-Truth images provided by the dataset are binarized images (8-bit), images with different channel numbers cannot be directly put into the discriminator for processing.
+As shown in Figure 2, except for the gray single-channel images, we process the remaining three single-channel images (red, green, and blue) using the Haar wavelet transform (one of the DWTs), and retain the LL subband images for normalization.
+Since the sizes of the retained LL subband images are $112\times 112$ (half of $224\times 224$), we resize the images to the original size of $224\times 224$ using the bicubic interpolation [37].
+In addition, as illustrated in Figure 2, the Ground-Truth images used for the four channels are different.
+For the gray channel, we directly use the Ground-Truth images provided by the dataset.
+For the remaining three channels (red, green, and blue), we employ the results of global binarization of the single-channel images processed by the DWT with normalization as the Ground-Truth images.
+The reasons for using DWT with normalization and their effects on the images are described in Section 4.5.
 3.2.1 Discrete Wavelet Transform
-In this work, we use Haar wavelet transform [33], which is one of the DWT, to process the input image.
-This process can be conceived as a bandpass filter, selectively allowing signals with frequencies akin to those of the wavelet basis function to traverse.
-The flowchart of this DWT processing is illustrated in Figure 2, which generates four subband images: one low-frequency subband image and three high-frequency subband images.
-This work takes the single-channel split images as input and processes the n-direction by low-pass filtering, as shown in the following equation:
+In this work, we use the Haar wavelet transform [38] (one of the DWTs) to process the input images.
+This processing can be equated to using a bandpass filter that selectively allows signals with frequencies similar to the wavelet basis functions to pass through.
+The flowchart of this processing is illustrated in Figure 3, which generates four subband images, such as the low-low (LL) subband images, the high-low (HL) subband images, the low-high (LH) subband images, and the high-high (HH) subband images.
+The low-pass filtered signal $V_{L}$ at position $(m,n)$ is obtained by convolving the input signal $x$ with the low-pass filter $g$, where the input signal is downsampled by a factor of 2, as shown in the following equation:
 $\begin{split}v_{L}[m,n]=\sum_{k=0}^{K-1}x[m,2n-k]g[k]\end{split}$ (2)
-Next, this work uses the low-pass filtering applied to the m-direction of Eq. (2) to obtain the following equation:
+where $x[m,2n-k]$ is the input signal $x$ at position $(m,2n-k)$, and this indexing suggests downsampling by a factor of 2.
+$g[k]$ is the filter coefficients for the low-pass filter, and $K$ is the length of the filter $g$.
+The low-pass filtered signal $V_{L}L$ at position $(m,n)$ is obtained by convolving the intermediate low-pass filtered signal $V_{L}$ (filtered along the columns) with the low-pass filter $g$, where the intermediate signal is downsampled by a factor of 2 along the row dimension, as shown in the following equation:
 $\begin{split}v_{LL}[m,n]=\sum_{k=0}^{K-1}v_{L}[2m-k,n]g[k]\end{split}$ (3)
-In contrast to the ground truth images provided by the dataset, the RGB single-channel images split from the original input images exhibit some substantial noise.
-This noise prevents them from being used directly in generating corresponding ground truth images through the generator and could potentially disrupt the performance of the trained adversarial network.
-DWT decomposes the image into both low-frequency and high-frequency information.
-The low-frequency component represents the mean value and the high-frequency component represents the difference value.
-The mean value, classified as low-frequency information, captures gradual changes and encodes both the picture’s contour information and approximate details.
-Conversely, the difference value, characterized as high-frequency information, captures rapid changes and encodes intricate picture details as well as local information (including noises).
-Therefore, we apply DWT and retain the LL subband images to input Stage-2, as shown in Figure 2, which effectively preserves contour information and reduces noise interferences.
-Figure 3: Results of Haar wavelet transform & normalization of RGB single-channel images split from the example image patches.
-Figure 4: Image normalization from image HW15 of DIBCO 2011 dataset, the above is an example of no normalization after DWT retains LL subband images, whereas the below is the example of normalization after DWT retains LL subband images.
+In contrast to the Ground-Truth images provided by the dataset, the three single-channel images (red, green, and blue) possess a significant amount of noise.
+The noise will prevent them from being used directly through the generator to generate the corresponding Ground-Truth images and possibly severely affect the performance of the trained GANs models.
+The DWT decomposes the images into both low-frequency and high-frequency information, where the low-frequency information represents the mean value, and the high-frequency information represents the difference value.
+The mean value captures gradients and encodes both contour information and proximate details of the images, and the difference value captures rapid changes and encodes complex image details and local information.
+Therefore, we apply the DWT and retain the LL subband images, as shown in Figure 3, which effectively preserves contour information and reduces the noise interference.
 3.2.2 Normalization
-Normalization entails constraining data within a desired range, with the goal of simplifying the data processing and expediting convergence. The proposed equation for this process is expressed as:
+Normalization entails constraining the data to the desired range with the aim of simplifying data processing and speeding up convergence. The proposed equation for this process is expressed as:
 $\begin{split}I_{N}=(\operatorname{IV_{nmax}}-\operatorname{IV_{nmin}})\frac{1}{1+e^{-\frac{I-\beta}{\alpha}}}+\operatorname{IV_{nmin}}\end{split}$ (4)
@@
 $\alpha$ defines the width of the input intensity range, and $\beta$ defines the range centered intensity.
-In the proposed method, we normalize the LL subband images to present initially incomparable data for comparison.
-The reason we choose the normalized LL subband image as the output is shown in Figure 3, where we apply DWT and normalization to each subband images.
-The illustrative example of the impact of normalization on RGB single-channel images after the DWT operation is provided in Figure 4, when applied to image HW15 from the DIBCO 2011 dataset, normalization effectively balances the RGB.
-In contrast, without applying normalization results in an output image, it heavily skews towards the blue component.
-In addition, we also show that the global binarization results of the images using normalization are more similar to the Ground Truth images provided by the dataset compared to the images only processed by DWT, and the experiments are explained in detail in Section 4.5.
-Figure 5: The architecture of Stage-2 of the proposed method.
-This method employs the U-Net++ [34, 35] architecture with EfficientNet-B6 [36] as the generator and improved PatchGAN [26] as the discriminator.
-The ground truth image during the training process is obtained from pixel-wise summation of the dataset provided image and the output from Stage-1.
-Figure 6: The architecture of Stage-3 of the proposed method.
-This stage comprises both local and global predictions.
-The local prediction involves training several $224\times 224$ image patches, while the global prediction involves directly resizing the original image to $512\times 512$ for training.
+In our method, we normalize the LL subband images by the DWT to present initially incomparable data for comparison.
+The reason we choose the normalized LL subband images as the output of Stage-1 is shown in Figure 4, where the foreground information in another subband images is not prominent compared to the LL subband images.
+Figure 5 illustrates the example of the effect of normalization on three single-channel patch images (red, green, and blue) processed by the DWT, which effectively balances three colors when applied to the “HW15” image of the DIBCO 2011 dataset [11].
+In contrast, the output images without normalization are heavily biased towards the blue component.
+In conclusion, the main difference between the results of Figure 5 (a) and (b) is that the images without normalization may lead to noise in the single-channel images, which seriously affect the pixel-level summation results.
+In addition, we also show that, compared to the binarization results of the images only processed by the DWT, the binarization results of the images using the DWT with normalization are more similar to the Ground-Truth images provided by the dataset.
+The related experiments are explained in detail in Section 4.5.
 3.3 Stage-2
 In Stage-2, the proposed method is illustrated in Figure 6.
-Initially, the original input images are split into three RGB single-channel images and one grayscale image.
-Each of the single-channel images is input into an independent generator, resulting in the use of four generators in Stage-2 of the proposed method.
-These four generators train the model separately, enabling this method to remove background information from the localized image blocks while extracting colored foreground details.
-Additionally, all single-channel images share the same discriminator to distinguish between the generated image and its corresponding ground truth image.
+Each single-channel patch image is input into an independent generator, so four generators are used in Stage-2 of the proposed method, which enable the method to extract color foreground information while removing background noise from the patch images.
+In addition, all single-channel images share the same discriminator to distinguish between the generated images and the corresponding Ground-Truth images.
 3.3.1 Loss functions of GAN
-To ensure a more robust convergence of the loss function, we follow the improved Wasserstein generative adversarial network with gradient penalty (WGAN-GP) [37] used by Suh et al.
-[31] applied to the objective function.
-Experiments by Bartusiak et al.
-[38] have demonstrated that the binary cross-entropy (BCE) loss yields superior performance compared to the $L1$ loss in binary classification tasks.
-Therefore, we chose the BCE loss instead of the $L1$ loss used in the Pix2Pix GAN [26].
-The loss function of the WGAN-GP, which incorporates the BCE loss, is formulated as follows:
+To ensure a more robust convergence of the loss, we follow an improved loss function of the Wasserstein generative adversarial network with gradient penalty (WGAN-GP) [39] used by Suh et al.
+[31], where WGAN-GP loss incorporates the BCE loss.
+Since Bartusiak et al.’s [40] experiments demonstrate that the binary cross-entropy (BCE) loss outperforms the $L1$ loss in binary classification tasks, Suh et al.’s [31] loss function employs the BCE loss instead of the $L1$ loss used in the Pix2Pix GAN [26].
+This improved loss function of the WGAN-GP is as follows:
+$\begin{split}\mathop{\mathbb{L}_{G}}=\mathop{\mathbb{E}_{x}}[D(G(x),x)]+\lambda\mathop{\mathbb{E}_{G(x),y}}[y\log G(x)\\ +(1-y)\log(1-G(x))]\end{split}$ (5)
 $\begin{split}\mathop{\mathbb{L}_{D}}=-\mathop{\mathbb{E}_{x,y}}[D(y,x)]+\mathop{\mathbb{E}_{x}}[D(G(x),x)]\\ +\alpha\mathop{\mathbb{E}_{x,\hat{y}\sim P_{\hat{y}}}}[(||\nabla_{\hat{y}}D(\hat{y},x)||_{2}-1)^{2}]\end{split}$ (6)
-$\begin{split}\mathop{\mathbb{L}_{G}}=\mathop{\mathbb{E}_{x}}[D(G(x),x)]+\lambda\mathop{\mathbb{E}_{G(x),y}}[y\log G(x)\\ +(1-y)\log(1-G(x))]\end{split}$ (6)
-where the penalty coefficient is $\alpha$, and the uniform sampling along a straight line between the ground truth distribution $P_{y}$ and the point pairs of the generated data distribution is $P_{\hat{y}}$.
+where the penalty coefficient is $\alpha$, and the uniform sampling along a straight line between the Ground-Truth distribution $P_{y}$, and the point pairs of the generated data distribution is $P_{\hat{y}}$.
 $\lambda$ is used to control the relative importance of different loss terms.
-In the discriminator, the generated image is distinguished from the real image by the loss function $\mathbb{L}_{D}$ in Eq.
+In the generator, the distances between the generated images and the Ground-Truth images in each color channel are minimized by the loss function $\mathbb{L}_{G}$ in Eq.
 (5).
-In the generator, the distance between the generated image and the ground truth image in each color channel is minimized by the loss function $\mathbb{L}_{G}$ in Eq.
+In the discriminator, the generated images are distinguished from the real images by the loss function $\mathbb{L}_{D}$ in Eq.
 (6).
-3.3.2 Network architecture
-Given the unpredictability of color document degradation, this work adopts four independent adversarial networks to extract text information from diverse colored backgrounds.
-This method mitigates color interference during document image binarization.
-For the network proposed in Stage-2, four independent generators apply the same encoder-decoder architecture.
-The encoder extracts features, while a shared decoder is employed by multiple encoders of varying depths, and subsequently the features are concatenated.
-A skip connection is used between the encoder and the decoders.
-The encoder facilitates downsampling and context extraction, while the decoders undertake upsampling and amalgamate both upscaled and downsampled features.
-Specifically, this work employs EfficientNet-B6 [36] as the encoder, which is a lightweight CNN known for its performance in image classification tasks.
-For the discriminator, we choose the one [26, 29, 30, 39, 40, 41] used by Suh et al.
-[31] due to its superior generalization capabilities.
-Additionally, instead of U-Net [21] used by Suh et al.
-[31], this work uses its improved model U-Net++ [34, 35] for semantic segmentation, which can further improve the model performance.
-The detailed procedure for training the proposed GAN models is illustrated in Algorithm 1.
-The training process is conducted in our training environment, which is described in Section 4.3, encompassing optimizer selection, learning rate configuration, and loss function definition.
-The algorithm demonstrates that the GAN model iteratively updates the loss function based on the authenticity assessment of the generated image by the discriminator.
-This iterative process enhances the ability of the generator to produce more real images.
 Algorithm 1 Training process for the proposed GAN models
@@
 21:  $\hat{y}\leftarrow[\hat{y}_{r},\hat{y}_{g},\hat{y}_{b}]$
+3.3.2 Network architecture
+For the network architecture proposed in Stage-2, the four independent generators use the same encoder-decoder architecture, where the encoder extracts the features and multiple encoders of different depths use a shared decoder and subsequently concatenate the features.
+A skip connection is used between the encoder and the decoders.
+The encoder facilitates downsampling and context extraction, while the decoders undertake upsampling and amalgamate both upscaled and downsampled features.
+Specifically, we employ EfficientNet-B6 [36] as the encoder, which is a lightweight convolutional neural network known for its performance in image classification tasks.
+Instead of U-Net [21] used by Suh et al.
+[31], we use the improved model U-Net++ [34, 35] as the network architecture for the generator, which can further improve the model performance.
+In addition, we follow the discriminator used by Suh et al. [31] as the discriminator in our network architecture due to its better generalization  [26, 29, 30, 41, 42, 43].
+The detailed procedure for training the proposed GANs models is illustrated in Algorithm 1.
+The training process is based on our constructed training environment, which includes the optimizer selection, learning rate, and loss function settings, as described in Section 4.3.
+The algorithm demonstrates that the GANs models iteratively update the loss function based on the discriminator’s assessment of the trueness of the generated images.
+This iterative process enhances the generator’s ability to generate more realistic images.
 3.4 Stage-3
-In Stage-3, a multi-scale adversarial network is employed to generate images for both local and global binarization results, enhancing the distinction between the background and text with higher accuracy.
-Contrary to Stage-2, which predominantly relies on local prediction using the image patches ($224\times 224$), Stage-3 conducts global prediction on the resized input image ($512\times 512$).
-This method reduces the contextual information loss in text caused by the local predictions of Stage-2.
+In Stage-3, we employ multi-scale GANs models to generate images for both local and global binarization results to enhance the distinction between foreground and background with higher accuracy.
+In contrast to Stage-2, which relies on the local prediction using patch images ($224\times 224$), Stage-3 performs the global prediction on the resized original input images ($512\times 512$).
+Therefore, this method reduces the contextual information loss in text caused by the local predictions of Stage-2.
 Figure 7 illustrates the process of two independent discriminators in Stage-3.
-Different from Stage-2 using 8-bit images for local prediction, Stage-3 uses 24-bit images as the input of the generator.
-Additionally, Figure 6 clarifies that the left-side input image comprises the image patches from the Stage-2 output, intended for local prediction.
-Conversely, the right-side input image represents the resized image from the original input for global prediction.
-In the binarization process of Stage-3, the network architecture of the generator remains consistent with Stage-2, differing only in the number of input channels.
+Different from Stage-2 using 8-bit images for the local prediction, Stage-3 uses 24-bit images as the inputs of the generator.
+In addition, Figure 7 illustrates that the left-side input images consist of the output images from Stage-2, which are used for the local prediction.
+In contrast, the right-side input images represent the resized images ($512\times 512$) from the original input images for the global prediction.
+It is worth noting that in the binarization process of Stage-3, the network architecture of the generator remains the same as in Stage-2, differing only in the number of input channels.
 4 Experiments
 4.1 Datasets
-4.1.1 DIBCO
-The Document Image Binarization Competition (DIBCO) provides several datasets: DIBCO 2009 [42], H-DIBCO 2010 [43], DIBCO 2011 [11], H-DIBCO 2012 [44], DIBCO 2013 [12], H-DIBCO 2014 [13], H-DIBCO 2016 [14], DIBCO 2017 [15], and H-DIBCO 2018 [16].
-These datasets comprise both grayscale and color images, including machine-printed as well as handwritten content.
+4.1.1 DIBCO Dataset
+The document image binarization competition (DIBCO) provides nine datasets [44, 45, 11, 46, 12, 13, 14, 15, 16]: DIBCO 2009, H-DIBCO 2010, DIBCO 2011, H-DIBCO 2012, DIBCO 2013, H-DIBCO 2014, H-DIBCO 2016, DIBCO 2017, and H-DIBCO 2018.
+These datasets comprise both gray-scale and color images, including machine-printed as well as handwritten content.
 Participants in the competition were tasked with designing algorithms capable of extracting binarized text images from both handwritten and machine-printed images within these datasets.
@@
 4.1.3 SMADI
-The Synchromedia Multispectral Ancient Document Images Dataset (SMADI) [1] consists of 240 multispectral images of 30 authentic historical handwritten letters.
-These letters, dating from the 17th to the 20th centuries, were written using iron gall ink.
+The synchromedia multispectral ancient document images (SMADI) dataset [1] consists of 240 multispectral images of 30 authentic historical handwritten letters.
+These letters, dating from the 17th to the 20th centuries, are written using iron gall ink.
 The original documents are deposited at the BAnQ Bibliotheque et Archives nationales du Quebec.
-A total of 240 images of real documents were meticulously captured, calibrated, and registered within this dataset.
-Each of these images was imaged with a CROMA CX MSI camera, resulting in 8 images per document.
+A total of 240 images of real documents are meticulously captured, calibrated, and registered within this dataset.
+Each of these images is imaged with a CROMA CX MSI camera, resulting in 8 images per document.
 4.1.4 Bickley Diary Dataset
-The Bickley Diary Dataset (BD) [46] has been generously donated to the Singapore Methodist Archives by Erin Bickley.
+The bickley diary (BD) dataset [48] has been generously donated to the Singapore Methodist Archives by Erin Bickley.
 This dataset comprises diaries in which the images have been affected by both light and fold damages, thereby complicating the process of identification.
 4.2 Evaluation Metric
-In comparison with other traditional and SOTA methods for document image binarization, based on the case that DIBCO datasets do not provide OCR output, we use the following five evaluation metrics:
+In comparison with other methods for document image enhancement and binarization, based on the case that DIBCO datasets do not provide OCR output, we use the following five metrics:
 4.2.1 F-measure (FM)
 ${\rm FM}=\frac{2\times{\rm Recall}\times{\rm Precision}}{{\rm Recall}+{\rm Precision}}$ (7)
-where the Recall is defined as $\frac{TP}{TP+FN}$; the Precision is defined as $\frac{TP}{TP+FP}$, and TP, FP, and FN represent the true positive, false positive, and false negative values, respectively.
+where the $Recall$ is defined as $\frac{TP}{TP+FN}$; the $Precision$ is defined as $\frac{TP}{TP+FP}$, and TP, FP, and FN represent the true positive, false positive, and false negative values.
 4.2.2 Pseudo-F-measure (p-FM)
 ${\rm p-FM}=\frac{2\times{\rm Recall_{ps}}\times{\rm Precision_{ps}}}{{\rm Recall_{ps}}+{\rm Precision_{ps}}}$ (8)
-where the Pseudo-Recall (${\rm Recall_{ps}}$) and the Pseudo-Precision (${\rm Precision_{ps}}$) both use the distance weights relative to the ground truth image profile to generate a new weighted ground truth image.
+where both pseudo-recall (${\rm Recall_{ps}}$) and pseudo-precision (${\rm Precision_{ps}}$) use distance weights relative to the contours of the Ground-Truth images to generate new weighted images.
 4.2.3 Peak signal-to-noise ratio (PSNR)
 ${\rm PSNR}=10\log(\frac{{\rm V}^{2}}{{\rm MSE}})$ (9)
-where ${\rm V}$ represents the difference between the foreground and background, and the Mean Square Error (MSE) is defined as:
+where ${\rm V}$ denotes the difference between the foreground and background, and the mean square error (MSE) is defined as:
 ${\rm MSE}=\frac{\sum_{x=0}^{M-1}\sum_{y=0}^{N-1}{[}L(x,y)-L^{\prime}(x,y){]}^{2}}{M\times N}$ (10)
@@
 4.2.4 Distance reciprocal distortion (DRD)
-${\rm DRD}$ represents a measure of visual distortion in the binary image, which is shown in the following equation:
+${\rm DRD}$ represents a measure of visual distortion in the binary images, which is shown in the following equation:
 ${\rm DRD}=\frac{{\rm\sum_{k}DRD_{k}}}{{\rm NUBN}}$ (11)
-where ${\rm NUBN}$ corresponds to the number of uneven $8\times 8$ blocks in the ground truth image, and ${\rm DRD_{k}}$ is defined as:
+where ${\rm NUBN}$ corresponds to the number of uneven $8\times 8$ blocks in the Ground-Truth images, and ${\rm DRD_{k}}$ is defined as:
 ${\rm DRD_{k}}=\sum_{i=-2}^{2}\sum_{j=-2}^{2}|{\rm G_{k}(i,j)-B_{k}(x,y)|\times N_{w}(i,j)}$ (12)
 where ${\rm DRD_{k}}$ indicates the distortion of the ${\rm k_{th}}$ flipped pixel, calculated using a $5\times 5$ normalized weight matrix ${\rm N_{w}}$.
-${\rm B_{k}}$ refers to the pixel of the binary image, and ${\rm G_{k}}$ denotes the pixel of the ground truth image.
+${\rm B_{k}}$ refers to the pixel of the binary images, and ${\rm G_{k}}$ denotes the pixel of the Ground-Truth images.
 4.2.5 Avg-Score (Avg)
-The Peak Signal-to-Noise Ratio (PSNR) provides a straightforward and intuitive quantitative metric for comparing the generated image with the ground truth image.
-Although it can be used as a supplementary metric, it should not be the only evaluation measure.
-The F-measure and p-F-measure are used to assess the accuracy of different methods, but a higher value in these metrics does not always correspond to a higher PSNR value.
-In our experiments, it is common to observe that a method achieves the SOTA PSNR value, while the F-measure and p-F-measure values are lower than those of other models.
-To solve this problem, we followed the Avg-Score metric used by Jemni et al.
+The PSNR metric provides a brief and intuitive quantitative way to compare the difference between the generated images and the Ground-Truth images, which can be used as a supplementary metric but should not be used as the only evaluation metric.
+The FM and p-FM metrics are used to assess the accuracy of different methods, but higher values of these metrics do not always correspond to higher PSNR values.
+Based on the results of our extensive experiments, it is common to observe that a method achieves the SOTA PSNR value, while the FM and p-FM values are lower than those of other models.
+To solve this problem, we follow the Avg-Score metric used by Jemni et al.
 [49] to compare the combined performance of different methods, as shown in the following equation:
 $Avg=\frac{FM+p\!-\!FM+PSNR+(100-DRD)}{4}$ (13)
-Table 1: Comparison of the PSNR values of the images from different training sets that have been processed using different methods with the corresponding ground truth images.
+Table 1: Quantitative comparison of PSNR values (dB) of images processed by different techniques with the corresponding Ground-Truth images on different training sets.
+The highest and 2nd highest PSNR values are in red and blue colors, respectively.
 Method DIBCO 2009 H-DIBCO 2010 H-DIBCO 2012 Bickley Diary PHIBD SMADI Mean Values
-Original Input 63.2879 61.3446 61.8542 57.6995 60.7355 61.2175 61.0232
-DWT 64.1154 71.4422 64.8151 53.7037 51.5094 60.1374 60.9539
-DWT&Norm 80.4286 86.1904 82.4625 74.2314 83.6403 78.5337 80.9145
-Table 2: Quantitative comparison (FM/p-FM/PSNR/DRD/Avg) of the models trained with different patch sizes ($224\times 224$/$256\times 256$) on multiple benchmark datasets.
-DIBCO 2011 DIBCO 2013 H-DIBCO 2014 H-DIBCO 2016 DIBCO 2017 DIBCO 2018 Mean Values
+Original 65.9812 62.7183 63.6814 59.3267 65.0543 62.5198 63.2136
+LL (DWT) 56.8869 62.9102 57.4896 50.4236 51.6242 57.1778 56.0854
+HL (DWT) 49.8898 49.7108 49.9529 50.8580 50.0277 49.7025 50.0236
+LH (DWT) 49.9061 49.7136 49.9578 50.8483 50.0311 49.6946 50.0253
+HH (DWT) 49.0228 48.8979 49.2434 50.0180 49.2218 48.9594 49.2272
+LL (DWT&Norm) 70.2888 77.0135 74.3312 65.1956 71.4992 69.4219 71.2917
+HL (DWT&Norm) 56.8078 58.2430 59.3027 56.6744 57.1121 56.7830 57.6497
+LH (DWT&Norm) 56.0911 58.7330 58.9662 57.0237 57.9381 56.9083 57.6101
+HH (DWT&Norm) 63.5701 64.5478 63.4654 61.8377 61.6075 61.7661 62.7991
+Table 2: Quantitative comparison (FM/p-FM/PSNR/DRD/Avg) of the proposed methods with different patch sizes ($224\times 224$ vs. $256\times 256$) on multiple benchmark datasets.
+DIBCO 2011 DIBCO 2013 H-DIBCO 2014 H-DIBCO 2016 DIBCO 2017 H-DIBCO 2018 Mean Values
 $256\times 256$ FM↑ 94.11 95.06 96.62 91.40 90.95 91.76 93.32
@@
 Avg↑ 77.64 77.95 79.05 76.38 75.34 77.00 77.22
-Figure 7: Visualization examples of the results of the proposed method at each stage.
-(a) the original input images, (b) the LL subband images obtained by Haar wavelet transform and normalization (Stage-1), (c) the enhanced images produced using our image enhancement method (Stage-2), (d) the binarized images obtained by combining local and global predictions (Stage-3), (e) the ground truth images.
-Table 3: Comparison of performance evaluation of the proposed method and other methods for document binarization on multiple benchmark datasets. Best and 2nd best performance are in red
-and blue colors, respectively. The data for the experimental results of other methods are obtained from [31, 48].
+Figure 8: Visualization examples of the results of the proposed method at each stage: (a) the original input images, (b) the LL subband images processed by Haar wavelet transform, (c) the LL subband images processed by Haar wavelet transform with normalization (Stage-1), (d) the enhanced images produced using our image enhancement method (Stage-2), (e) the binarized images obtained by combining local and global predictions (Stage-3), (f) the Ground-Truth images.
+Table 3: Quantitative comparison of performance evaluation (FM/p-FM/PSNR/DRD/Avg) of the proposed method and other methods for document image binarization on multiple benchmark datasets.
+Best and 2nd best performance are in red and blue colors, respectively.
+The data for the experimental results of other methods are obtained from [31, 50].
 Method FM↑ p-FM↑ PSNR↑ DRD↓ Avg↑
@@
 Ours 92.89 96.96 20.39 2.23 77.00
-(f) DIBCO 2018
-Figure 8: Example of document image binarization results from input image HW1 of DIBCO 2011 dataset by different methods: (a) original input, (b) Otsu [6], (c) Niblack [5], (d) Sauvola [7], (e) Vo [23], (f) He [24], (g) Suh [31], (h) Ours, (i) ground truth.
-Figure 9: Example of document image binarization results from input image HW5 of DIBCO 2013 dataset by different methods: (a) original input, (b) Otsu [6], (c) Niblack [5], (d) Sauvola [7], (e) Vo [23], (f) He [24], (g) Suh [31], (h) Ours, (i) ground truth.
+(f) H-DIBCO 2018
+Figure 9: Example of document image binarization results from the input image “HW1” of DIBCO 2011 dataset by different methods: (a) original input image, (b) Otsu’s [6], (c) Niblack’s [5], (d) Sauvola et al.’s [7], (e) Vo et al.’s [23], (f) He et al.’s [24], (g) Suh et al.’s [31], (h) Ours, and (i) Ground-Truth image.
+Figure 10: Example of document image binarization results from input image “HW5” of DIBCO 2013 dataset by different methods: (a) original input image, (b) Otsu’s [6], (c) Niblack’s [5], (d) Sauvola et al.’s [7], (e) Vo et al.’s [23], (f) He et al.’s [24], (g) Suh et al.’s [31], (h) Ours, and (i) Ground-Truth image.
 4.3 Implementation Details
 4.3.1 Data preparation
-The performance of a neural network depends on the quality and quantity of its training data.
-In general, the more training data enhances the better the performance of the neural network.
-To compare the performance of the proposed method with other methods, we follow the strategy described in [23, 24, 31] for constructing the training and test sets.
+The model performance of the neural network depends on the quality and quantity of its training data.
+In general, the more training data improves the better model performance.
+To fairly compare the performance of the proposed method with other methods, we follow the strategy described in [23, 24, 31] to construct the training and test sets.
 Specifically, the training set includes a total of 143 images from DIBCO 2009 [44], H-DIBCO 2010 [45], H-DIBCO 2012 [46], PHIBD [47], SMADI [1], and BD [48] dataset.
-The test sets comprise a total of 82 images from DIBCO 2011 [11], DIBCO 2013 [12], H-DIBCO 2014 [13], H-DIBCO 2016 [14], DIBCO 2017 [15], and H-DIBCO 2018 [16].
-To ensure a fair comparison of performance among all methods, we used the same data augmentation for our method and other SOTA methods.
-During the training process of the local prediction network, this work splits the input images into patches with $224\times 224$ in size.
-These patches were sampled with scale factors of ${0.75,1.25,1.5}$, subsequently resized to $224\times 224$, and rotated by 270°.
-In contrast, for training the global prediction network, this work employed scaling without rotation for data augmentation, using a scaling size of $512\times 512$.
-To diversify the training data, rotation augmentation was applied to the global prediction, involving rotations of 90°, 180°, and 270°, as well as horizontal and vertical flips.
+While the test sets comprise a total of 82 images from DIBCO 2011 [11], DIBCO 2013 [12], H-DIBCO 2014 [13], H-DIBCO 2016 [14], DIBCO 2017 [15], and H-DIBCO 2018 [16].
+To ensure a fair comparison of the performance of all methods, for the proposed method and other methods, we apply the same data augmentation techniques to the training data.
+In the training process of the GANs models of Stage-2, we split the original input images into patches of $224\times 224$ size, and these patches are sampled with scale factors of ${0.75,1,1.25,1.5}$.
+In addition, we rotate the image by $270\degree$ to obtain more training data.
+In contrast, for training GAN model on resized $512\times 512$ images of Stage-3, we employ scaling without rotation for data augmentation.
+To extend the training data, we employ the rotation augmentation technique during the training process of this model, including $90\degree$, $180\degree$ and $270\degree$ rotations, horizontal and vertical flips.
 4.3.2 Training
-The experiments were conducted using the PyTorch framework in Python. We use NVIDIA RTX 3090 GPUs for all the following experiments.
-In this study, we employ the generator comprising EfficientNet-B6 [36] with UNet++ [34, 35].
-Regarding the training hyperparameters, Stage-2 and Stage-3 share similar settings, differing primarily in the number of training epochs: 150 epochs for global prediction and 10 epochs for the remaining networks.
-The choice of optimizer was Adam [49], with a learning rate set at $2\times 10^{-4}$.
-Additionally the generator employed $\beta_{1}=0.5$, while the discriminator used $\beta_{2}=0.999$.
+The experiments are conducted using the PyTorch framework in Python. We use single NVIDIA RTX 4090 GPU for all the experiments.
+In this work, we employ the generator consisting of EfficientNet-B6 [36] with U-Net++ [34, 35].
+About the training hyper-parameters, the settings of Stage 2 and Stage 3 are similar, with the only difference in the number of training epochs.
+Specifically, for the network of the global prediction we set it to 150 epochs and for the remaining networks we set it to 10 epochs.
+We select the Adam [51] optimizer, with an initial learning rate set to $2\times 10^{-4}$.
+In addition, we set $\beta_{1}=0.5$ for the generator, and $\beta_{2}=0.999$ for the discriminator.
 4.3.3 Pre-training
-To enhance training efficacy due to limited data availability, we employed EfficientNet-B6 [36] as the encoder in the generator, utilizing weights pretrained on the ImageNet dataset [50].
+Due to the limited data available, to improve the training efficiency, we choose EfficientNet-B6 [36] as the encoder of the generator, utilizing the pre-trained weights on the ImageNet dataset [52].
 4.4 Comparison with other Methods
-In our previous conference paper [51], we compared various deep learning-based methods and found that some of these methods [10, 28, 29] employed a “leave-one-out” strategy for selecting the training set.
-For instance, when the DIBCO 2013 dataset is used as the test set, all other datasets are utilized as the training set.
-Although our method still demonstrates superior performance at the SOTA level in such case [51], we consider it unfair to compare them directly because our model uses a smaller training set.
-Therefore, this work excludes these methods from this paper and only retains those that use the exact same training set for comparison.
+In our previous conference paper [53], we compared different deep learning-based methods and found that some of these methods [10, 28, 29] employ “leave-one-out” strategy for constructing the training set.
+For instance, when the DIBCO 2013 [12] dataset is used as the test set, all other DIBCO datasets are used as the training set.
+Although our method still shows excellent performance in such case of [53], we consider it is unfair to compare them directly because our model uses a much smaller training set.
+Therefore, we exclude these methods from the methods we compare in this paper, and retain only those methods [5, 6, 7, 17, 18, 23, 24, 31] that use the same training set for comparison.
 4.5 Image Enhancement Experiment
-In Stage-1, we applied DWT and normalization to the original input images.
-While mathematical theory [52] suggests that these processed images can retain contour information and reduce noise more efficiently, it is important to demonstrate their impact on experimental results.
-The previous conference paper [51] on this work presented ablation experiments on DWT and normalization.
-Based on this, it is better to intuitively show that, compared to the original input images, the DWT and normalization processed images are closer to the ground truth images.
-In Stage-1 of this work, we set the threshold to 0.3 to binarize different single-channel images to obtain corresponding binarized images.
-In the experiment of this section, we applied the above image binarization to the original input images, the images processed with DWT, and the images processed with both DWT and normalization.
-Subsequently, we calculated the PSNR values of the binarized images compared to the ground truth images using these three methods, and the results are presented in Table 2.
-The PSNR mean values of the original input images are 61.0232.
-In contrast, the PSNR mean values of the images after DWT processing are slightly lower at 60.9539.
-However, the PSNR mean values of the images processed with both DWT and normalization are 80.9145, which means they are closer to the ground truth images.
-This demonstrates that it is more reasonable to choose the DWT and normalization processed image rather than the original input image and the DWT processed image.
+The previous conference paper [53] of this work presents ablation study on the DWT with normalization.
+Based on this, it is better to intuitively show that, compared to the binarization results of the original input images, the binarization results of the images processed by the DWT with normalization are closer to the Ground-Truth images.
+While the mathematical theory [54] suggests that the images processed by the Haar wavelet transform can retain contour information and reduce noise more efficiently, it is important to demonstrate their impact on experimental results.
+In Stage-1 of this work, we set the threshold to 0.5 to binarize different single-channel $224\times 224$ patch images (red, green, and blue) to obtain the corresponding binarization results.
+In the experiment of this section, we apply the above image binarization to the original input images, the images only processed by the DWT, and the images processed by the DWT with normalization.
+Subsequently, we calculate the PSNR values of the binarized images by different techniques compared to the Ground-Truth images, and the results are presented in Table 1.
+Compared with the PSNR mean value of the original input images of 63.2136dB, the PSNR mean values of the four subband images (LL, HL, LH, HH) only processed by the DWT are all lower, which are 56.0854dB, 50.0236dB, 50.0253dB, and 49.2272dB, respectively.
+In addition, the PSNR mean values of the three subband images (HL, LH, LL) processed by the DWT with normalization are lower than that of the original input images, which are 57.6497dB, 57.6101dB, and 62.7991dB, respectively, while the PSNR mean value of the LL subband images processed by the DWT with normalization reaches 71.2917dB, which means they are closer to the Ground-Truth images.
+The results of this experiment show that the choice of the DWT with normalization is more reasonable than only the DWT for processing the images.
 4.6 Comparison Experiment
-To ensure fairness in the experiments, we used the same training set and data augmentation techniques described in Section 4.3.1 for all the compared models.
+To ensure fairness in the experiments, we use the same training set and data augmentation techniques described in Section 4.3.1 for all the compared models.
 Unlike Suh et al.
-[31], who split the input images into $256\times 256$ patches, our proposed method uses $224\times 224$ patches.
-To demonstrate the positive impact of this change on the model performance, we conducted experiments comparing models using $224\times 224$ patches with those using $256\times 256$ patches, keeping all other settings unchanged.
+[31], who split the input images into $256\times 256$ patches, our proposed method sets the patch size to $224\times 224$.
+To demonstrate the positive impact of this change on the model performance, we conduct the experiments comparing models using $224\times 224$ patches with those using $256\times 256$ patches, keeping all other settings unchanged.
 Table 2 presents the performance of the models across different patch sizes on six DIBCO datasets.
-As established in Section 4.5, the images processed with DWT and normalization outperform those processed with either the original input images or only with DWT.
-Consequently, all models in this section utilize images processed with DWT and normalization.
+Since we have demonstrated in Section 4.5 that the DWT with normalization technique outperforms the only DWT technique, the models in Table 2 all use the DWT with normalization technique.
 According to Table 2, the mean values of FM, p-FM, PSNR, DRD, and Avg for the model based on $224\times 224$ patches are 93.57, 96.83, 20.59, 2.09, and 77.22, respectively, which are superior to those for the model based on $256\times 256$ patches.
-This indicates that the model based on $224\times 224$ patches performs better, and therefore we will use this model for subsequent comparisons.
+This indicates that the model based on $224\times 224$ patches has better model performance, and therefore we will use the model based on $224\times 224$ patches for subsequent comparisons.
 4.7 Experimental Results
-Different from the conference paper [51] of this work, which uses option 3 to train the model, the proposed method in this paper applies DWT and normalization to the images in Stage-1 as input to train the model.
-The performance of the trained model will then be compared with that of other methods.
-Figure 7 presents the visualization results at each stage of the trained model.
-Specifically, five images were randomly selected from the PHIBD and Bickley Diary datasets to demonstrate the step-by-step process of image enhancement and binarization using the proposed method.
-Figure 7 (b) represents the result of retaining the LL subband images after applying DWT and normalization.
-This stage performs noise reduction on the original input image, which is used as the ground truth image for the generator.
-Figure 7 (c) shows the result of image enhancement achieved using adversarial networks, where the image has the background color removed and the color of the text highlighted.
-Figure 7 (d) is the final output image obtained using the proposed whole method.
-Notably, it is evident that our final output result closely approximates Figure 7 (e), the ground truth image.
-Table 3 presents the evaluation results of different models on six DIBCO datasets.
-Overall, the proposed method achieves the top two best performance in terms of FM, p-FM, PSNR, and DRD, with its Avg value reaching the SOTA level across all six datasets.
-Notably, the performance of the proposed method on the DIBCO 2011 and 2013 datasets is particularly good, obtaining the best values for each of the evaluated metrics.
+Different from the conference paper [53] of this work that uses the DWT with normalization to only process the input images, this paper uses the DWT with normalization to process both the input images and the Ground-Truth images of the three single-channel images (red, green, and blue) in Stage-1.
+Figure 8 presents the visualization results at each stage of the proposed model.
+Specifically, four images are randomly selected from the PHIBD [47] and BD [48] datasets to demonstrate the step-by-step process of document image enhancement and binarization.
+Figure 8 (b) represents the results of the LL subband images procossed by the DWT, and (c) represents the results of the LL subband images procossed by the DWT with normalization.
+It can be seen that there is a significant difference between (b) and (c), which performs noise reduction on the original input images.
+Figure 8 (d) shows the effect of image enhancement achieved using the GANs models, where the images remove the background color and highlight the color of the text.
+Figure 8 (e) presents the final output images obtained using the whole proposed method.
+It is worth noting that the final outputs are very close to Figure 8 (f) the Ground-Truth images.
+Table 9(f) presents the evaluation results of different models on six DIBCO datasets.
+Overall, the proposed method achieves the top two best performance in terms of FM, p-FM, PSNR, and DRD, while its Avg value reaching the SOTA level across all six datasets.
+Notably, the performances of the proposed method on the DIBCO 2011 [11] and H-DIBCO 2018 [16] datasets are particularly good, obtaining the best values for each of the evaluation metrics.
 For instance, on the DIBCO 2011 dataset, the p-FM value of the proposed method is 97.47, significantly surpassing the second best value of 95.93.
-Similarly, on the DIBCO 2018 dataset, the FM value of the proposed method is 92.89, considerably higher than the second best value of 91.86.
-These results indicate that the proposed method is more effective than other methods for the binarization of degraded color documents.
-In addition to assessing method performance by comparing quantitative metrics across various methods, this work aims to provide readers with an intuitive understanding of the distinctions by the visualization results.
-Consequently, we have chosen images from the DIBCO 2011 and 2013 datasets to illustrate examples of binarization results.
-Figures 8 and 9 show the binarization results of images HW1 and HW5 by different methods.
-These figures visually demonstrate that the proposed method outperforms deep learning based methods and the traditional binarization methods in terms of the shadow and noise elimination.
-Furthermore, the proposed method is good at preserving textual content while effectively mitigating the presence of the shadows and noise.
-5 Conclusion
-This paper introduces a novel three-stage method based on DWT and GANs to apply the task of text degradation in degraded document images with color backgrounds.
-The proposed method involves preserving the LL subband images through DWT to eliminate image noise, serving as the ground truth images for the generator.
-It trains independent networks utilizing RGB three single-channel images and one grayscale image, followed by a fusion of local and global prediction networks for binarization.
-As far as our knowledge extends, this work first represents the three-stage method that combines traditional image processing with deep learning for document binarization.
-The experimental results demonstrate that employing U-Net++ with EfficientNet-B6 as the generator and PatchGAN as the discriminator outperforms performance compared to other SOTA methods.
+Similarly, on the H-DIBCO 2018 dataset, the FM value of the proposed method is 92.89, considerably higher than the second best value of 91.86.
+These results indicate that the proposed method is more effective than other methods using the same training set for document image enhancement and binarization.
+In addition to assessing method performance by comparing the quantitative metrics of different methods, this work aims to allow the reader to intuitively understand the differences between the methods by visualizing the binarization results.
+Figures 9 and 10 show the examples the binarization results by different methods for the image “HW1” of the DIBCO 2011 dataset, and “HW5” of the DIBCO 2013 dataset.
+These figures visually demonstrate that the proposed method outperforms the traditional threshold methods and deep learning-based methods in terms of the shadow and noise elimination.
+In addition, the proposed method retains the text content well while effectively mitigating the effects of shadows and noise.
+Specifically, as shown in Figure 10 (g) and (h), the proposed method is more efficient in eliminating noise compared to Suh et al.’s [31] method.
+5 Conclusion and Future Work
+This paper introduces an effective three-stage network method for enhancement and binarization of gray-scale and color degraded document images.
+Compared to other methods, the proposed method is more efficient in removing noise from degraded document images, and this effect applies to both gray-scale and color images.
+As far as our knowledge extends, this work first represents the three-stage network method that combines traditional image processing with deep learning for document image enhancement and binarization.
+The experimental results show that the performance of the proposed method using U-Net++ with EfficientNet-B6 as generator and PatchGAN as discriminator is better compared to other methods.
+In addition, based on the experimental results, this paper also demonstrates that for document images, the binarization results of the images processed by the DWT with normalization are closer to the Ground-Truth images than the binarization results of the original input images.
+We believe that this paper is of great significance to the reconstruction and analysis of ancient documents.
+In future work, we will apply the proposed method to more degraded document images from different cultures to extract useful information to help culture organizations to analyze the history of ancient civilizations.
+In particular, we will focus on the complex structure and polysemous characters, such as Chinese, Egyptian, Arabic, etc.
 6 Declarations
@@
 6.3 Declaration of Generative AI and AI-assisted technologies in the writing process
-The authors only use generative artificial intelligence (AI) and AI-assisted technologies to improve language.
+The authors only use generative artificial intelligence (AI) and AI-assisted technologies to improve the language.